The third grade classroom has a perimeter of 130 feet. If it is 25 feet wide,how many feet long is the classroom.

Answer by macston(5194)

L=length; W=width; P=perimeter
P=2(L+W) Divide each side by 2
P/2=L+W Subtract W from each side
(P/2)-W=L Plug in numbers
(130/2)-25=L
65-25=L
40=L ANSWER The length of the classroom is 40 feet.